At a distance of 73 cm, a mosquito can barely be heard at a sound level of 4.1 dB.
What is the intensity of this sound?


If the same mosquito were flying at a distance of 1.3 cm from your ear, what would be the intensity of its sound?


If there were two mosquitoes 1.3 cm from your ear, how loud would they sound? (dB).


How loud would the mosquito sound at 1.3 cm from your ear (dB)?

Hint: sound intensities add together. Decibels do not.
Answer #1

-12 (a) 1b1 4,1 dB = 10 los 7 ole I = 2.57*10th Wim - Answer = Chevy SE 2-576161317) - 8.1 x 109 Wilma pa lo log 212 = 10 lo